# GAEA: A Geolocation Aware Conversational Model

### v2

@@ 6 Conclusion
Figure 9: Geo-localization qualitative example. GAEA’s performance on geo-localization tasks is compared to open-source LMMs using CityGuessr, DollarStreet, and GeoDE datasets.
Figure 10: SVQA qualitative example. GAEA’s performance on SVQA tasks is compared to open-source and proprietary LMMs.
We organize the rest of the Supplementary Material as follows: In Section 7, we provide the additional details of our dataset, GAEA-1.6M . In Section 8, we provide additional baseline results on GAEA. Section Section 9 discusses the reproducibility of the GAEA-1.6M and provides details on privacy, safety, and broader impact.
@@ 7 Addendum to the Dataset
Figure 11: These two images display examples of what we consider as easy questions. Easy questions include the questions that pertain to easily identifiable landmarks that are associated with celebrated locations.
Figure 12: The two images above denote examples of what we consider as hard questions. Hard questions include the questions that prompt the model to answer specific details pertaining to locations.
Figure 13: Through our explanatory captions (LVQA), we introduce the reasoning capabilities in our GAEAto identify the geographical information about that visual sample. Outlined in bold, we provide explanations of the correlation between specific visual cues and their associated geographical contexts, encouraging GAEAto refine its reasoning capabilities.
@@
Total Questions 1,580,531
Total geo-localization questions 822,951
Total explanatory captions 384,947
Total open-ended questions 267,668
Total multiple-choice questions 48,673
Total true/false questions 56,292
@@
7.1 Challenges with Open Street Maps (OSM)
OpenStreetMaps (OSM) [38] is a rich data source for geospatial applications. It contains a wide variety of geographic and infrastructure-related information. Using such a vast open-source dataset, we can collect data about stationary objects in the world, including infrastructure, topological information, various types of amenities (e.g., schools, hospitals, restaurants), transportation networks, international country boundaries, historical and cultural sides, and natural features (e.g., forests, rivers, and seas). Each feature from the OSM dataset has several associated features, such as names and physical characteristics. In GAEA-1.6M, we geocode the visual sample with its GPS coordinates and use the location information (longitude and latitude) as a query to the OSM database to fetch geospatial information in 1 KM radius and further utilize that information to generate question-answer pairs for the training of GAEA.
7.2 Statistics
GAEA-1.6M covers 234 different countries and territories, and 41,481 cities. Table 3 denotes the exact sample numbers.
7.3 Motivation for constructing GAEA-1.6M
The dataset contains a wide spectrum of questions, with varying difficulty levels. Figure 11 outlines a category consisting of easily geo-localizable landmarks such as the Statue of Liberty, and Figure 12 represents one of the many difficult questions found in our dataset, which attempts to fine-tune an LMM to respond with location information that an average human could struggle to identify [36]. In Figure 13, we show a full example of one of our LVQA prompts.
Training an LMM with some of the metadata found in OSM [38] can be a challenging task. For instance, questions like, What are the hours of operation of the nearest coffee shop located in this area? could be difficult for a model to learn effectively and respond to accurately. We encourage the research community to explore and develop methods that could help LMMs meaningfully represent such fine-grained geo-localization information.
@@ 8.1 Baselines
Additionally, we compared the performance of GAEA against six state-of-the-art (SoTA) geo-localization models, namely PlaNet [53], CPlaNet [45], ISNs [25], TransLocator [52], GeoDecoder [18], and PIGEON [23] on three standard geo-localization benchmarks including IM2GPS [24], IM2GPS3k [51], and GWS15k [18]. For a comprehensive overview of the hyperparameter configuration for GAEA, see Table 4 in the Supplementary Material.
Figure 14: MCQ qualitative example. GAEA’s performance on MCQ answering tasks is compared to open-source and proprietary LMMs.
Figure 15: T/F qualitative example. GAEA’s performance on T/F answering tasks is compared to open-source and proprietary LMMs.
@@ 8.2 Prompts Used During Training and Inference
Evaluation Prompt for LVQA Evaluate the following predicted answer by comparing it to the provided ground truth. Focus on the accuracy of 1) location prediction, 2) cultural aspect matching, 3) consistency and quality of reasoning, 4) specificity and relevance, 5) and fluency and clarity. Question: {question} Ground Truth: {ground_truth} Model Prediction: {predicted_answer} Instructions: How accurately does the predicted answer identify the specific country, city, or state mentioned in the ground truth? Does the predicted answer capture and reflect the cultural aspects present in the ground truth? Is the predicted answer logically consistent and demonstrates sound reasoning based on the information provided? Does the predicted answer provide specific information that is directly relevant to the question and closely aligns with the ground truth? Is the language in the predicted answer fluent, clear, and well-articulated? Provide a single overall score out of 10, based on these five criteria, weighing the criteria in the order listed, with location relevance and cultural aspect matching receiving the most weight. Return only the numeric score, without additional commentary.
Figure 16: Evaluation prompt for longer form open-ended questions (LVQA). We assess model predictions based on location prediction, cultural aspect matching, and quality of reasoning, with an emphasis on location relevance.
Figure 17: We showcase the performance of various LMMs on four diverse question types. GAEA outperforms on average across all question forms. GPT-4o achieves the highest accuracy on long questions.
Figure 18: Task-specific prompts used to train and evaluate GAEA
Figure 19: Example of the country-specific clues we used to generate reasoning questions.
Evaluation Prompt for SVQA Evaluate the following predicted answer by comparing it to the provided ground truth. Focus on the accuracy of 1) location prediction, and 2) specificity and relevance. Question: {question} Ground Truth: {ground_truth} Model Prediction: {predicted_answer} Scoring Guidelines: High score: Predicted response closely matches the specific location and provides specific information that closely aligns with the ground truth. Low score: Predicted response lacks knowledge or is unrelated to the ground truth Provide a score out of 10 for each criterion. Return only the numeric score, without additional commentary
Figure 20: Evaluation prompt for free-form open-ended questions (SVQA). We assess model predictions based on location accuracy, specificity, and correctness.
Evaluation Prompt for MCQ/TF Evaluate the following answer based on Accuracy: Question: {question} Ground Truth: {ground_truth} Model Prediction: {predicted_answer} Instructions: Match the meaning of the ground truth with the model prediction. If it matches, give a score of 10. Otherwise, give a score of 0. Strictly return only the numeric score, without any additional commentary.
Figure 21: Evaluation prompt for multiple-choice (MCQ) and true/false (TF) questions.
@@ 8.3 Prompts Used in Evaluation
To obtain our results from Figure Figure 17, we use three prompts in evaluating the questions in GAEA-Bench. Since the question types of our benchmark vary, a multi-prompt approach is needed. We have reformatted certain elements of the actual prompts for conciseness, though figures Figures 16, 20 and 21 closely reflect these three prompts.
@@
Model data type bfloat16
Maximum context length of LLM 128,000
Attention type flash attention
@@ 8.4 Training Hyperparameters
We perform single-stage training on our baseline [12] using GAEA-1.6M. The training is conducted for 1 epoch with a global batch size of 128, utilizing gradient accumulation steps of 4 to optimize resource usage for small batch sizes. The initial learning rate is set to $10^{-5}$, using a cosine learning rate scheduler to provide a smooth decay in the learning rate for effective convergence. A weight decay of 0 is applied to avoid penalizing weights during updates, which can be advantageous for certain model architectures. The warmup ratio is configured at 0.03 to ensure a gradual increase in the learning rate during the initial training phase, stabilizing early optimization. We employed low-Rank adaptation (LoRA) [26] for efficient fine-tuning, with a rank, $r=16$, $\alpha=32$, and a dropout rate of 0.01, enabling targeted model adjustments with minimal overhead. The model operates in bfloat16 precision to balance computational efficiency and numerical stability. A maximum context length of 128,000 allows for processing extremely long sequences, while flash attention enhances the computational efficiency of attention mechanisms, especially for extended contexts. These settings collectively optimize the model’s performance and adaptability for vision-language tasks. We list the training hyperparameters in the Table 4.
8.5 Additional Qualitative Results
In this Section, we discuss additional qualitative results of GAEA and compare them with selected open-source and proprietary models (as mentioned in Table 1 in the main paper). Figure 9 presents a comparison of city-country predictions against other competing models. We also show the qualitative results of GAEA on short version questions (SVQA), multiple-choice questions (MCQs), and true or false questions (TF) in Figures 10, 14 and 15 respectively. For these Figures, we highlight correct predictions with green, while incorrect predictions are marked as red. Together with our quantitative metrics, we conclude that GAEA appears to outperform many of these models across various tasks.
9 Reproducibility, Privacy, Safety, and Broader Impact
GAEA-1.6M takes the first step in infusing conversational elements into the geo-localization task. The dataset is open-source, and we plan to release it via an academic website for research, academic, and commercial use. The dataset is protected under the CC-BY license of Creative Commons, which allows the users to distribute, remix, adapt, and build upon the material in any medium or format, as long as the creator is attributed. The license allows GAEA-1.6M for commercial use. As the authors of this manuscript and collectors of this dataset, we reserve the right to distribute the data. Additionally, we provide the code, data, and instructions needed to reproduce the main experimental baseline results, and the statistics pertinent to the dataset. We specify all the training details (e.g., data splits, hyperparameters, model-specific implementation details, compute resources used, etc.).
The dataset can be used by multiple domain experts. Its application includes but is not only limited to tourist assistance, government analysts, and GeoGuessr [2] enthusiasts. Although we do not find any foreseeable harm that the dataset can pose to human society, it is always possible that some individual or an organization can use this idea to devise a technique that can appear harmful to society and can have evil consequences. However, as authors, we are absolutely against any detrimental usage of this dataset, regardless of whether it is by an individual or an organization, under profit or non-profitable motivation, and we pledge not to support any detrimental endeavors concerning our data or the idea therein.

### v3

@@ 7 Acknowledgement
We organize the Supplementary Material as follows: In Section 8, we provide additional details of our dataset, GAEA-1.4M . In Section 9, we summarize baselines we compared against, describe prompts used during training, inference and evaluation, and provide training details and additional results on GAEA.
Figure 9: Examples of country-specific geographical clues used in our dataset. These clues are used to support geographically grounded reasoning in our QA generation.
@@ 8 Addendum to the Dataset
Figure 10: Examples of the four question types in our dataset: SVQA, MCQ, TF, and LVQA. Each type targets a distinct reasoning skill grounded in geographical, visual, or contextual understanding. Our dataset has three categories, including Geolocalization, Reasoning(LVQA), and Conversational (SVQA, MCQ, TF) QAs, as shown in the figure.
8.1 GeoGuessr Context Clues
To support geographically-grounded reasoning in our dataset, we incorporate country-specific context clues inspired by the popular game GeoGuessr [2]. We web-crawled and extracted 129 high-quality clues from Plonkit [4], a community-driven open-source resource widely used by over 65 million GeoGuessr players. To improve country coverage and address missing entries (e.g., France, New Zealand), we augmented the set with 58 additional GPT-4o-generated clues, carefully aligned with Plonkit’s descriptive style. This expanded our coverage to 187 countries. These clues highlight distinctive, often visual, geographical features, such as Bangladesh’s unique turquoise license plates, Australia’s widespread gum trees, or the ice-covered terrain of Grise Fiord in Canada (see Figure 9). These examples are leveraged to generate reasoning-based multiple-choice and open-ended questions. Furthermore, we enrich each sample with auxiliary metadata: Köppen-Geiger climate classification [13], traffic orientation from WorldStandards [7], land cover data from EarthEnv [1], and scene labels derived from the Places2 database [62].
8.2 Challenges with Open Street Maps (OSM)
OpenStreetMaps (OSM) [40] is a rich data source for geospatial applications. It contains a wide variety of geographic and infrastructure-related information. Using such a vast open-source dataset, we can collect data about stationary objects in the world, including infrastructure, topological information, various types of amenities (e.g., schools, hospitals, restaurants), transportation networks, international country boundaries, historical and cultural sides, and natural features (e.g., forests, rivers, and seas). Each feature from the OSM dataset has several associated features, such as names and physical characteristics. In GAEA-1.4M, we geocode the visual sample with its GPS coordinates and use the location information (longitude and latitude) as a query to the OSM database to fetch geospatial information within a  1 KM radius and further utilize that information to generate question-answer pairs for the training of GAEA.
8.3 Training Examples and Data Statistics
GAEA-1.4M covers 234 different countries and territories, and 41,481 cities. Table 3 shows the statistics of GAEA-1.4M in detail. Figure 10 shows some qualitative examples of various question types in our GAEA-1.4M  leveraging OSM, and GeoGuessr geographical clues for constructing conversational QA pairs.
@@ 8.3 Training Examples and Data Statistics
Total Questions 1,432,519
Total Geo-Loc QAs 822,951
Total LVQAs 236,935
Total SVQAs 267,668
Total MCQs 48,673
Total True/False QAs 56,292
@@
Figure 11: Task-specific training prompts used to instruct the model for each of the five question types in GAEA. Each prompt is carefully designed to elicit a targeted form of geographical reasoning. These prompts ensure consistent and interpretable outputs during both the training and evaluation phases while training GAEA.
@@ 9.1 Baselines
Additionally, we compared the performance of GAEA against six state-of-the-art (SoTA) specialized geolocalization models, namely PlaNet [56], CPlaNet [47], ISNs [27], TransLocator [54], GeoDecoder [18], and PIGEON [25] and open-source geolocalization LMMs GeoReasoner [34] and GaGA [20] on three standard geo-localization benchmarks including IM2GPS [26], IM2GPS3k [53], and GWS15k [18]. We also compare our city-country classification performance with other LMMs on 3 benchmarks CityGuessr68k [30], GeoDE [43] and DollarStreet [23]. Preprocessing for these benchmarks is described in Section 5.1 in the main paper.
Figure 12: Evaluation prompts used for GPT-based scoring of model predictions across SVQA, LVQA, MCQ, and TF question types in the GAEA-Bench. Each prompt specifies detailed criteria, such as location accuracy, cultural relevance, specificity, and fluency, used to assign a numeric score for qualitative assessment.
@@ 9.3 Prompts Used in Evaluation
Figure 12 presents the task-specific prompts used for evaluating model-generated answers via GPT-based assessment. For SVQA and LVQA, the prompts emphasize multi-faceted criteria such as location accuracy, specificity, cultural relevance, and reasoning quality, encouraging a nuanced evaluation of open-ended responses. In contrast, MCQ and TF tasks are scored based on strict binary accuracy by matching the predicted answer to the ground truth. These structured prompts ensure consistent, interpretable, and criterion-aligned evaluation across different QA types.
@@
Precision bfloat16
@@ 9.4 Training Hyperparameters
We perform single-stage training on the baseline, Qwen2.5-VL [12] using GAEA-1.4M. The training is conducted for 1 epoch with a global batch size of 128, accumulating gradients every 4 steps. The initial learning rate is set to $10^{-5}$, using a cosine learning rate scheduler to provide a smooth decay in the learning rate. The warmup ratio is configured at 0.03. We performed LoRA[28]-finetuning with a rank, $r=16$, $\alpha=32$, and a dropout rate of 0.01. The model operates in bfloat16 precision. We also use flash attention [19]. We list the training hyperparameters in the Table 4.
Figure 13: Qualitative Geolocalization comparison showing GAEA’s ability to predict accurate locations across varied types of scenes and landscapes around the world.
Figure 14: Qualitative SVQA comparison showing GAEA’s ability to provide accurate, location-specific answers where other LMMs fail.
Figure 15: Qualitative MCQs comparison showing GAEA’s ability to provide accurate answers where other LMMs fail.
Figure 16: Qualitative True/False comparison showing GAEA’s ability to provide accurate answers where other LMMs fail.
Figure 17: We showcase the performance of various LMMs on four diverse question types. GAEA outperforms on average across all question forms.
9.5 Additional Results
In this Section, we discuss additional qualitative results of GAEA and compare them with selected open-source and proprietary models (as mentioned in Table 1 in the main paper). Figure 13 presents a comparison of city-country predictions against other competing models. We also show the qualitative results of GAEA on short questions (SVQA), multiple-choice questions (MCQs), and true or false questions (TF) in Figures 14, 15 and 16. For these Figures, we highlight correct predictions with green, while incorrect predictions are marked as red. Quantitative results on GAEA-Bench are summarized in Figure 17.